(d) Assume that 4 years from now you will need $1,000. Your bank compounds interest at an 8% annual rate.
(i) How much must you deposit 1 year from now to have a balance of Sh.100, 000 at Year 4? [2 marks]
(ii) If you want to make equal payments at the end of Years 1 through 4 to accumulate the Sh.100, 000, how large must each of the 4 payments be?
1
Expert's answer
2020-12-21T15:25:18-0500

I) Amount, "A=P(1+r\/100)^{n}"

"100000=P(1\n+8\/100)^{4}=1.36049P"

"P=100000\/1.36049= sh. 73502.93"


II) "100000-73502.93=sh.26,497.07"

each of the four payments "=26497.07\/4=sh. 6,624.27"